A 5.580 kg block of wood rests on a steel desk. The coefficient of static friction between the block and the desk is ?s=0.455 and the coefficient of kinetic friction is ?k=0.255. At time ?=0, a force ?=15.3 N is applied horizontally to the block. State the force of friction applied to the block by the table at times ?=0 and ?>0.

?=0

N?>0

NConsider the same situation, but this time the external force ? is 30.9 N. Again, state the force of friction acting on the block at times ?=0 and ?>0.?=0

N?>0

N
Answer #1

ا E =1m Solution A ms S.58oleg F. = 15:39 Ms = 0.455 + Na mg = 5.58oX9.8 N- 54.684 maxinum static friction = tmax = MSN trax